# Quasi-boolean groups

### v2

@@ 7. Quillen’s theorems and their consequences
Quasi-canonical homomorphisms $k\in H^{1}(G,\mathbb{Z}/2)$ can be characterised by the following property: for every $n>0$ and $c\in H^{n}(G,\mathbb{Z}/2)$ we have $c^{2}=c\cup k^{n}$.
@@
An elementary $p$-group $H$ is a group isomorphic to $(\mathbb{Z}/p)^{n}$ for some $n$. The rank of $H$ is $n$, that is, its dimension as a vector space over $\mathbb{Z}/p$. The elementary rank of a pro-$p$ group $G$ is the supremum of all natural numbers $r$ such that $G$ has a subgroup isomorphic to an elementary $p$-group of rank $r$.
@@
    every involution $x\in G$ is not in the kernel of any quasi-canonical homomorphism,
@@
To prove $(i)$, let $x\in G$ be an involution, let $H\subseteq G$ be the subgroup generated by $x$ and let $i\colon H\to G$ be the inclusion map. Assume that there is a quasi-canonical homomorphism $k\in H^{1}(G,\mathbb{Z}/2)$ whose restriction to $H$ is zero, or equivalently $i^{*}(k)=0$. By Corollary 7.3, there is an $n>0$ and a $c\in H^{n}(G,\mathbb{Z}/2)$ such that $i^{*}(c)\in H^{n}(H,\mathbb{Z}/2)$ is non-zero. Then
$00\neq i^{*}(c)^{2}=i^{*}(c^{2})=i^{*}(c\cup k^{n})=i^{*}(c)\cup i^{*}(k)^{n}=$ 0
using that $H^{\bullet}(H,\mathbb{Z}/2)$ is isomorphic to a polynomial ring in one variable over $\mathbb{Z}/2$, but this is a contradiction. Therefore $(i)$ holds.
To prove $(ii)$, assume to the contrary that $G$ contains a subgroup $H$ isomorphic to an elementary $2$-group of rank $2$. Then the kernel of the restriction of a quasi-canonical homomorphism of $H$ is non-trivial. This contradicts part $(i)$, so claim $(ii)$ is true. ∎
@@ 2. $(ii)$
Since graded Boolean algebras have no nilpotent elements, we get that $f$ is injective by condition $(i)$ of Definition 7.12. Next we show that $f$ is surjective. Since $f$ is clearly an isomorphism in degree zero, it will be sufficient to show that for every positive integer $n$ and $x\in C^{n}$ there is a $y\in B^{n}$ such that $f(y)=x$. By assumption, there is a positive integer $m$ and a $z\in B^{nm}$ such that $f(z)=x^{m}$. Because $B^{\bullet}$ is a graded Boolean algebra the $m$-th power map $B^{n}\to B^{nm}$ is an isomorphism, so there is a $y\in B^{n}$ such that $y^{m}=z$, and hence $f(y)^{m}=f(y^{m})=f(z)=x^{m}$. Since $C^{\bullet}$ is also a graded Boolean algebra the $m$-th power map $C^{n}\to C^{nm}$ is an isomorphism, so $f(y)=x$. ∎
Definition 7.14.
For every profinite group $G$, let $\mathfrak{A}(G)$ denote the set of all subgroups of $G$ which are finite elementary abelian $p$-groups. Note that all such subgroups are closed, since $G$ is Hausdorff. Also note that $\mathfrak{A}(G)$ form a category where morphisms are maps $f\colon A\to B$ such that there is an $x\in G$ such that $f(y)=x^{-1}gx$ for every $y\in A$. Note that the assignment $G\mapsto\mathfrak{A}(G)$ is functorial, that is for every continuous homomorphism $h:G\to H$ there is an induced functor $\mathfrak{A}(h)\colon\mathfrak{A}(G)\to\mathfrak{A}(H)$. For every open normal subgroup $N\triangleleft G$, let $\mathfrak{A}(G,N)$ denote the image of $\mathfrak{A}(G)$ under the functor $\mathfrak{A}(\pi_{N})\colon\mathfrak{A}(G)\to\mathfrak{A}(G/N)$ induced by the quotient map $\pi_{N}:G\to G/N$.
Definition 7.15.
Let $\mathbf{DGAC}_{p}$ denote the category of graded anti-commutative algebras over $\mathbb{Z}/p$. For every profinite group $G$, let $F_{G}\colon\mathfrak{A}(G)\to\mathbf{DGAC}_{p}$ be the functor given by the rule $A\to H^{\bullet}(A,\mathbb{Z}/p)$. For every open normal subgroup $N\triangleleft G$, let $F_{G,N}\colon\mathfrak{A}(G,N)\to\mathbf{DGAC}_{p}$ denote the restriction of $F_{G/N}$ onto $\mathfrak{A}(G,N)$. For every $G$ and $N$ as above, let $\underline{H^{\bullet}}_{\mathfrak{A}}(G,\mathbb{Z}/p)$ and $\underline{H^{\bullet}}_{\mathfrak{A}}(G,N,\mathbb{Z}/p)$ denote the inductive limit of the functor $F_{G}$ and $F_{G,N}$, respectively.
Definition 7.16.
Let $G$ and $N$ be as above. Then for every $A\in\mathfrak{A}(G,N)$, let $\pi_{A,N}\colon A\to\pi_{N}(A)$ be the map induced by the restriction of $\pi_{N}$ onto $A$. Note that for every
$\underline{c}=\{c_{A}\in H^{\bullet}(A,\mathbb{Z}/p)\mid A\in\mathfrak{A}(G,N)\}\in\underline{H^{\bullet}}_{\mathfrak{A}}(G,N,\mathbb{Z}/p)$
@@
$\mathfrak{a}_{N}(\underline{c})=\{\pi_{N,A}^{*}(c_{\pi_{N}(A)})\in H^{\bullet}(A,\mathbb{Z}/p)\mid A\in\mathfrak{A}(G)\}$
lies in $\underline{H^{\bullet}}_{\mathfrak{A}}(G,\mathbb{Z}/p)$ and the map $\mathfrak{a}_{N}\colon\underline{H^{\bullet}}_{\mathfrak{A}}(G,N,\mathbb{Z}/p)\to\underline{H^{\bullet}}_{\mathfrak{A}}(G,\mathbb{Z}/p)$ is a homomorphism of graded anti-commutative algebras over $\mathbb{Z}/p$. Let $H^{n}_{\mathfrak{A}}(G,\mathbb{Z}/p)$ denote the union of the images of these homomorphisms as $N$ ranges over the set of all open normal subgroups of $G$. Since these images form an inductive system, $H^{n}_{\mathfrak{A}}(G,\mathbb{Z}/p)$ is a graded subalgebra of $\underline{H^{\bullet}}_{\mathfrak{A}}(G,\mathbb{Z}/p)$. Clearly, when $G$ is finite, we have $H^{\bullet}_{\mathfrak{A}}(G,\mathbb{Z}/p)=\underline{H^{\bullet}}_{\mathfrak{A}}(G,\mathbb{Z}/p)$.
@@
Let $G$ be a profinite group of elementary rank at most $1$ such that there is an open normal subgroup of $G$ without $p$-torsion. Then $H^{n}_{\mathfrak{A}}(G,\mathbb{Z}/p)=C(\mathcal{X}_{p}^{*}(G),\mathbb{Z}/p)$ for every $n>0$.
@@
Let $F(\mathcal{X}_{p}^{*}(G),\mathbb{Z}/p)$ denote the group of all $\mathbb{Z}/p$-valued functions on $\mathcal{X}_{p}^{*}(G)$. Since for every non-zero $A\in\mathfrak{A}(G)$, we have $A\cong\mathbb{Z}/p$. Hence $H^{n}(A,\mathbb{Z}/p)=\mathbb{Z}/p$, every element $c\in\underline{H}^{n}_{\mathfrak{A}}(G,\mathbb{Z}/p)$ gives rise to a function $\mathcal{Y}_{p}^{*}(G)\to\mathbb{Z}/p$ which is conjugation-invariant, and hence descends to a function $f(c)\colon\mathcal{X}_{p}^{*}(G)\to\mathbb{Z}/p$. The map $f\colon\underline{H}^{n}_{\mathfrak{A}}(G,\mathbb{Z}/p)\to F(\mathcal{X}_{p}^{*}(G),\mathbb{Z}/p)$ is an isomorphism. So the precise meaning of our claim is that the image of $H^{n}_{\mathfrak{A}}(G,\mathbb{Z}/p)$ under $f$ and $C(\mathcal{X}_{p}^{*}(G),\mathbb{Z}/p)$ are equal as subgroups of $F(\mathcal{X}_{p}^{*}(G),\mathbb{Z}/p)$.
First let $c\in H^{n}_{\mathfrak{A}}(G,\mathbb{Z}/p)$ be arbitrary. Then there is an open normal subgroup $N\triangleleft G$ and a $d\in\underline{H}^{n}_{\mathfrak{A}}(G,N,\mathbb{Z}/p)$ such that $c=\mathfrak{a}_{N}(d)$. Since for every pair $M,N$ of open, normal subgroups of $G$ such that $M\subseteq N$ the image of $\mathfrak{a}_{M}$ contains the image of $\mathfrak{a}_{N}$, we may assume that $N$ does not contain $p$-torsion without the loss of generality by shrinking $N$ if it is necessary. Then $\pi_{N}$ induces a map $\pi_{N}^{\#}\colon\mathcal{X}_{p}^{*}(G)\to\mathcal{X}_{p}^{*}(G/N)$, and $f(c)$ is the composition of $\pi_{N}^{\#}$ with the function $\mathrm{Im}(\pi_{N}^{\#})\to\mathbb{Z}/p$ corresponding to $d$. Since both $\pi_{N}^{\#}$ and the latter function are continuous, we get that $f(c)$ is continuous, too.
Now let $c\in C(\mathcal{X}_{p}^{*}(G),\mathbb{Z}/p)$ be arbitrary, and let $g\colon\mathcal{Y}_{p}^{*}(G)\to\mathbb{Z}/p$ be the continuous function we get by composing the quotient map $\mathcal{Y}_{p}^{*}(G)\to\mathcal{X}_{p}^{*}(G)$ with $c$. Since the translates of normal open subgroups of $G$ form a sub-basis for its topology, for every $x\in\mathcal{Y}_{p}^{*}(G)$ there is a normal open subgroup $N_{x}\triangleleft G$ such that $g$ is constant on $xN_{x}\cap\mathcal{Y}_{p}^{*}(G)$. Since there is an open normal subgroup of $G$ without $p$-torsion, the subspace $\mathcal{Y}_{p}^{*}(G)$ is closed, and hence compact, so there is a finite subset $S\subseteq\mathcal{Y}_{p}^{*}(G)$ such that $\mathcal{Y}_{p}^{*}(G)\subseteq\bigcup_{x\in S}xN_{x}$.
Since $N=\bigcap_{x\in S}N_{x}$ is the intersection of finitely many normal open subgroups, it is a normal open subgroup, too. We may even assume that $N$ does not contain $p$-torsion without the loss of generality by shrinking $N$ if it is necessary, as above. Then $g$ is constant on $xN\cap\mathcal{Y}_{p}^{*}(G)$ for every $x\in\mathcal{Y}_{p}^{*}(G)$, and hence $c$ is the composition of $\pi_{N}^{\#}$ above with a function $d\colon\mathrm{Im}(\pi_{N}^{\#})\to\mathbb{Z}/p$. ∎
Definition 7.18.
Let $G$ be as above, and, for every $A\in\mathfrak{A}(G)$, let $i_{A}\colon A\to G$ be the inclusion map. For every $c\in H^{\bullet}(G,\mathbb{Z}/p)$, the collection
$q_{G}(c)=\{i_{A}^{*}(c)\in H^{\bullet}(A,\mathbb{Z}/p)\mid A\in\mathfrak{A}(G)\}$
lies in $\underline{H}^{\bullet}_{\mathfrak{A}}(G,\mathbb{Z}/p)$, since every conjugation of $G$ induces the identity on $H^{\bullet}(G,\mathbb{Z}/p)$. The resulting map $q_{G}\colon H^{\bullet}(G,\mathbb{Z}/p)\to\underline{H}^{\bullet}_{\mathfrak{A}}(G,\mathbb{Z}/p)$ is a homomorphism of graded anti-commutative algebras over $\mathbb{Z}/p$.
Lemma 7.19.
The homomorphism $q_{G}$ maps $H^{\bullet}(G,\mathbb{Z}/p)$ into $H^{\bullet}_{\mathfrak{A}}(G,\mathbb{Z}/p)$.
Proof.
Let $c\in H^{n}(G,\mathbb{Z}/p)$ be arbitrary. Then there is an open normal subgroup $N\triangleleft G$ and a $d\in H^{n}(G/N,\mathbb{Z}/p)$ such that $c=\pi^{*}_{N}(d)$. Let $\phi_{N}\colon\underline{H}^{n}_{\mathfrak{A}}(G/N,\mathbb{Z}/p)\to\underline{H}^{n}_{\mathfrak{A}}(G,N,\mathbb{Z}/p)$ be the map induced by the inclusion functor $\mathfrak{A}(G,N)\to\mathfrak{A}(G/N)$. Clearly $q_{G}(c)=\mathfrak{a}_{N}(\phi_{N}(q_{G/N}(d)))$, and hence the claim holds. ∎
@@
$q_{G}\colon H^{\bullet}(G,\mathbb{Z}/p)\to H^{\bullet}_{\mathfrak{A}}(G,\mathbb{Z}/p)$
@@
This theorem was proved by Quillen when $G$ is finite (see [13, Theorem 7.1 on page 567] which is actually a more general result). At the end of [11] (see 8.6 and 8.7 on pages 279–80) Scheiderer pointed out that there is an easy limit argument to derive the theorem above as a corollary. In fact, he proved that $q_{G}$ satisfies property $(i)$ in Definition 7.12. We complete his argument by showing $(ii)$ for the convenience of the reader.
Definition 7.21.
For every pair $M,N$ of open, normal subgroups of $G$ such that $M\subseteq N$ let $\pi_{M,N}\colon G/M\to G/N$ be the quotient map, and let $\mathfrak{A}(G,M,N)$ denote the image of the functor $\mathfrak{A}(\pi_{M,N})\colon\mathfrak{A}(G/M)\to\mathfrak{A}(G/N)$.
@@
For every open, normal subgroup $N\triangleleft G$ there is open, normal subgroup $M\triangleleft G$ such that $M\subseteq N$ and $\mathfrak{A}(G,M,N)=\mathfrak{A}(G,N)$.

### v3

@@ 7. Quillen’s theorems and their consequences
Quasi-canonical homomorphisms $k\in H^{1}(G,\mathbb{Z}/2\mathbb{Z})$ can be characterised by the following property: for every $n>0$ and $c\in H^{n}(G,\mathbb{Z}/2\mathbb{Z})$, we have $c^{2}=c\cup k^{n}$. To prove this assertion, we write $H^{\bullet}=H^{\bullet}(G,\mathbb{Z}/2\mathbb{Z})$ as the connected sum of $D^{\bullet}$ and $B^{\bullet}$. Then $k\in H^{1}(G,\mathbb{Z}/2\mathbb{Z})$ is quasi-canonical if and only if its image under the quotient map $H^{\bullet}/D^{\bullet>0}\to B^{\bullet}$ is the identity map since $D^{\bullet>0}$ is the nilradical of $H^{\bullet}$ as pointed out in the proof of Lemma 7.7. The identity $c^{2}=c\cup k^{n}$ holds for all elements of $B^{\bullet}$, and this identity is equivalent to $k$ being the identity in $B$ by definition of the multiplication in $B^{\bullet}$ in Definition 7.4. The same identity trivially holds for all $c$ in $D^{\bullet>0}$ and any $k$.
@@
An elementary $p$-group $H$ is a group isomorphic to $(\mathbb{Z}/p\mathbb{Z})^{n}$ for some $n$. The rank of $H$ is $n$, that is, its dimension as a vector space over $\mathbb{Z}/p\mathbb{Z}$. The elementary rank of a pro-$p$ group $G$ is the supremum of all natural numbers $r$ such that $G$ has a subgroup isomorphic to an elementary $p$-group of rank $r$.
@@
    every involution $x\in G$ is not in the kernel of any quasi-canonical homomorphism;
@@
To prove $(i)$, let $x\in G$ be an involution, let $H\subseteq G$ be the subgroup generated by $x$ and let $i\colon H\to G$ be the inclusion map. Let $i^{\bullet}\colon H^{\bullet}(G,\mathbb{Z}/2\mathbb{Z})\to H^{\bullet}(H,\mathbb{Z}/2\mathbb{Z})$ denote the pullback homomorphism on cohomology. Assume that there is a quasi-canonical homomorphism $k\in H^{1}(G,\mathbb{Z}/2\mathbb{Z})$ whose restriction to $H$ is zero, or equivalently $i^{\bullet}(k)=0$. By Corollary 7.3, there is an $n>0$ and a $c\in H^{n}(G,\mathbb{Z}/2\mathbb{Z})$ such that $i^{\bullet}(c)\in H^{n}(H,\mathbb{Z}/2\mathbb{Z})$ is non-zero. Then, by Remark 7.9,
$0\neq i^{\bullet}(c)^{2}=i^{\bullet}(c^{2})=i^{\bullet}(c\cup k^{n})=i^{\bullet}(c)\cup i^{\bullet}(k)^{n}=0$
using that $H^{\bullet}(H,\mathbb{Z}/2\mathbb{Z})$ is isomorphic to a polynomial ring in one variable over $\mathbb{Z}/2\mathbb{Z}$, but this is a contradiction. Therefore $(i)$ holds.
To prove $(ii)$, assume to the contrary that $G$ contains a subgroup $H$ isomorphic to an elementary $2$-group of rank $2$. Then the kernel of the restriction of a quasi-canonical homomorphism to $H$ is non-trivial. This contradicts part $(i)$, so claim $(ii)$ is true. ∎
@@ 2. $(ii)$
Since graded Boolean algebras have no nilpotent elements, we get that $f$ is injective by condition $(i)$ of Definition 7.12. Next we show that $f$ is surjective. Since $f$ is clearly an isomorphism in degree zero, it will be sufficient to show that for every positive integer $n$ and $x\in C^{n}$ there is a $y\in B^{n}$ such that $f(y)=x$. By condition $(ii)$ of Definition 7.12, there is a positive integer $m$ and a $z\in B^{nm}$ such that $f(z)=x^{m}$. Since $B^{\bullet}$ is a graded Boolean algebra, the $m$-th power map $B^{n}\to B^{nm}$ is an isomorphism, so there is a $y\in B^{n}$ such that $y^{m}=z$, and hence $f(y)^{m}=f(y^{m})=f(z)=x^{m}$. Since $C^{\bullet}$ is also a graded Boolean algebra, the $m$-th power map $C^{n}\to C^{nm}$ is an isomorphism, so $f(y)=x$. ∎
Notation 7.14.
For every profinite group $G$, let $\mathfrak{A}(G)$ denote the set of all subgroups of $G$ which are finite elementary abelian $p$-groups. We note that, since $G$ is Hausdorff, all such subgroups are closed by [2, I §8.2, Proposition 4 on page 77]. Also note that $\mathfrak{A}(G)$ form a category where morphisms are maps $f\colon A\to B$ such that there is an $x\in G$ such that $f(y)=x^{-1}yx$ for every $y\in A$. Note that the assignment $G\mapsto\mathfrak{A}(G)$ is functorial, that is, for every continuous homomorphism $h\colon G\to H$ there is an induced functor $\mathfrak{A}(h)\colon\mathfrak{A}(G)\to\mathfrak{A}(H)$. For every open normal subgroup $N\triangleleft G$, let $\mathfrak{A}(G,N)$ denote the image of $\mathfrak{A}(G)$ under the functor $\mathfrak{A}(\pi_{N})\colon\mathfrak{A}(G)\to\mathfrak{A}(G/N)$ induced by the quotient map $\pi_{N}\colon G\to G/N$.
Notation 7.15.
Let $\mathbf{DGAC}_{p}$ denote the category of graded anti-commutative algebras over $\mathbb{Z}/p\mathbb{Z}$. For every profinite group $G$, let $F_{G}\colon\mathfrak{A}(G)\to\mathbf{DGAC}_{p}$ be the functor given by the rule $A\mapsto H^{\bullet}(A,\mathbb{Z}/p\mathbb{Z})$. For every open normal subgroup $N\triangleleft G$, let $F_{G,N}\colon\mathfrak{A}(G,N)\to\mathbf{DGAC}_{p}$ denote the restriction of $F_{G/N}$ onto $\mathfrak{A}(G,N)$. For every $G$ and $N$ as above, let $\underline{H}^{\bullet}_{\mathfrak{A}}(G,\mathbb{Z}/p\mathbb{Z})$ and $\underline{H}^{\bullet}_{\mathfrak{A}}(G,N,\mathbb{Z}/p\mathbb{Z})$ denote the inductive limit of the functors $F_{G}$ and $F_{G,N}$, respectively.
Notation 7.16.
Let $G$ and $N$ be as above. Then for every $A\in\mathfrak{A}(G,N)$, let $\pi_{N,A}\colon A\to\pi_{N}(A)$ be the map induced by the restriction of $\pi_{N}$ onto $A$. We let $\pi_{N,A}^{\bullet}\colon H^{\bullet}(\pi_{N}(A),\mathbb{Z}/p\mathbb{Z})\to H^{\bullet}(A,\mathbb{Z}/p\mathbb{Z})$ denote the pullback homomorphism on cohomology. Note that for every
$\underline{c}=\{c_{A}\in H^{\bullet}(A,\mathbb{Z}/p\mathbb{Z})\mid A\in\mathfrak{A}(G,N)\}\in\underline{H}^{\bullet}_{\mathfrak{A}}(G,N,\mathbb{Z}/p\mathbb{Z})$
@@
$\mathfrak{a}_{N}(\underline{c})=\{\pi_{N,A}^{\bullet}(c_{\pi_{N}(A)})\in H^{\bullet}(A,\mathbb{Z}/p\mathbb{Z})\mid A\in\mathfrak{A}(G)\}$
lies in $\underline{H}^{\bullet}{\mathfrak{A}}(G,\mathbb{Z}/p\mathbb{Z})$ and the map $\mathfrak{a}_{N}\colon\underline{H}^{\bullet}_{\mathfrak{A}}(G,N,\mathbb{Z}/p\mathbb{Z})\to\underline{H}^{\bullet}_{\mathfrak{A}}(G,\mathbb{Z}/p\mathbb{Z})$ is a homomorphism of graded anti-commutative algebras over $\mathbb{Z}/p\mathbb{Z}$. Let $H^{\bullet}_{\mathfrak{A}}(G,\mathbb{Z}/p)$ denote the union of the images of these homomorphisms as $N$ ranges over the set of all open normal subgroups of $G$. Since these images form an inductive system, $H^{\bullet}_{\mathfrak{A}}(G,\mathbb{Z}/p\mathbb{Z})$ is a graded subalgebra of $\underline{H}^{\bullet}_{\mathfrak{A}}(G,\mathbb{Z}/p\mathbb{Z})$. Clearly, when $G$ is finite, we have $H^{\bullet}_{\mathfrak{A}}(G,\mathbb{Z}/p\mathbb{Z})=\underline{H}^{\bullet}_{\mathfrak{A}}(G,\mathbb{Z}/p\mathbb{Z})$.
@@
Let $G$ be a profinite group of elementary rank at most $1$ such that there is an open normal subgroup of $G$ without $p$-torsion. Then $H^{n}_{\mathfrak{A}}(G,\mathbb{Z}/p\mathbb{Z})=C(\mathcal{X}_{p}^{*}(G),\mathbb{Z}/p\mathbb{Z})$ for every $n>0$ where $C(\mathcal{X}_{p}^{*}(G),\mathbb{Z}/p\mathbb{Z})$ denotes the ring of continuous functions from $\mathcal{X}_{p}^{*}(G)$ to $\mathbb{Z}/p\mathbb{Z}$, where the latter is equipped with the discrete topology.
Remark 7.18.
Let $F(\mathcal{X}_{p}^{*}(G),\mathbb{Z}/p\mathbb{Z})$ denote the group of all $\mathbb{Z}/p\mathbb{Z}$-valued functions on $\mathcal{X}_{p}^{*}(G)$. For every non-zero $A\in\mathfrak{A}(G)$, we have $A\cong\mathbb{Z}/p\mathbb{Z}$. Hence $H^{n}(A,\mathbb{Z}/p\mathbb{Z})=\mathbb{Z}/p\mathbb{Z}$, every element $c\in\underline{H}^{n}_{\mathfrak{A}}(G,\mathbb{Z}/p\mathbb{Z})$ gives rise to a function $\mathcal{Y}_{p}^{*}(G)\to\mathbb{Z}/p\mathbb{Z}$ which is conjugation-invariant, and hence descends to a function $f(c)\colon\mathcal{X}_{p}^{*}(G)\to\mathbb{Z}/p\mathbb{Z}$. The map $f\colon\underline{H}^{n}_{\mathfrak{A}}(G,\mathbb{Z}/p\mathbb{Z})\to F(\mathcal{X}_{p}^{*}(G),\mathbb{Z}/p\mathbb{Z})$ is an isomorphism. So the precise meaning of the claim of Proposition 7.17 is that the image of $H^{n}_{\mathfrak{A}}(G,\mathbb{Z}/p\mathbb{Z})$ under $f$ and $C(\mathcal{X}_{p}^{*}(G),\mathbb{Z}/p\mathbb{Z})$ are equal as subgroups of $F(\mathcal{X}_{p}^{*}(G),\mathbb{Z}/p\mathbb{Z})$.
Proof of Proposition 7.17.
First let $c\in H^{n}_{\mathfrak{A}}(G,\mathbb{Z}/p\mathbb{Z})$ be arbitrary. Then there is an open normal subgroup $N\triangleleft G$ and a $d\in\underline{H}^{n}_{\mathfrak{A}}(G,N,\mathbb{Z}/p\mathbb{Z})$ such that $c=\mathfrak{a}_{N}(d)$. Since for every pair $M,N$ of open normal subgroups of $G$ such that $M\subseteq N$ the image of $\mathfrak{a}_{M}$ contains the image of $\mathfrak{a}_{N}$, we may assume that $N$ does not contain $p$-torsion without the loss of generality by shrinking $N$ if it is necessary. Then $\pi_{N}$ induces a map $\pi_{N}^{\#}\colon\mathcal{X}_{p}^{*}(G)\to\mathcal{X}_{p}^{*}(G/N)$, and $f(c)$ is the composition of $\pi_{N}^{\#}$ with the function $\mathrm{Im}(\pi_{N}^{\#})\to\mathbb{Z}/p\mathbb{Z}$ corresponding to $d$. Since both $\pi_{N}^{\#}$ and the latter function are continuous, we get that $f(c)$ is continuous, too.
Now let $c\in C(\mathcal{X}_{p}^{*}(G),\mathbb{Z}/p\mathbb{Z})$ be arbitrary, and let $g\colon\mathcal{Y}_{p}^{*}(G)\to\mathbb{Z}/p\mathbb{Z}$ be the continuous function we get by composing the quotient map $\mathcal{Y}_{p}^{*}(G)\to\mathcal{X}_{p}^{*}(G)$ with $c$. Since the translates of normal open subgroups of $G$ form a sub-basis for the topology on $G$, for every $x\in\mathcal{Y}_{p}^{*}(G)$ there is an open normal subgroup $N_{x}\triangleleft G$ such that $g$ is constant on $xN_{x}\cap\mathcal{Y}_{p}^{*}(G)$. Since there is an open normal subgroup of $G$ without $p$-torsion, the subspace $\mathcal{Y}_{p}^{*}(G)$ is closed, and hence compact, so there is a finite subset $S\subseteq\mathcal{Y}_{p}^{*}(G)$ such that $\mathcal{Y}_{p}^{*}(G)\subseteq\bigcup_{x\in S}xN_{x}$.
Since $N=\bigcap_{x\in S}N_{x}$ is the intersection of finitely many normal open subgroups, it is an open normal subgroup, too. We may even assume that $N$ does not contain $p$-torsion without the loss of generality by shrinking $N$ if it is necessary, as above. Then $g$ is constant on $xN\cap\mathcal{Y}_{p}^{*}(G)$ for every $x\in\mathcal{Y}_{p}^{*}(G)$, and hence $c$ is the composition of $\pi_{N}^{\#}$ above with a function $d\colon\mathrm{Im}(\pi_{N}^{\#})\to\mathbb{Z}/p\mathbb{Z}$. ∎
Notation 7.19.
Let $G$ be as above, and, for every $A\in\mathfrak{A}(G)$, let $i_{A}\colon A\to G$ be the inclusion map. For every $c\in H^{\bullet}(G,\mathbb{Z}/p\mathbb{Z})$, the collection
$q_{G}(c)=\{i_{A}^{\bullet}(c)\in H^{\bullet}(A,\mathbb{Z}/p\mathbb{Z})\mid A\in\mathfrak{A}(G)\}$
lies in $\underline{H}^{\bullet}_{\mathfrak{A}}(G,\mathbb{Z}/p\mathbb{Z})$ since every conjugation of $G$ induces the identity on $H^{\bullet}(G,\mathbb{Z}/p\mathbb{Z})$. The resulting map $q_{G}\colon H^{\bullet}(G,\mathbb{Z}/p\mathbb{Z})\to\underline{H}^{\bullet}_{\mathfrak{A}}(G,\mathbb{Z}/p\mathbb{Z})$ is a homomorphism of graded anti-commutative algebras over $\mathbb{Z}/p\mathbb{Z}$.
Lemma 7.20.
The homomorphism $q_{G}$ maps $H^{\bullet}(G,\mathbb{Z}/p\mathbb{Z})$ into $H^{\bullet}_{\mathfrak{A}}(G,\mathbb{Z}/p\mathbb{Z})$.
@@
Let $c\in H^{n}(G,\mathbb{Z}/p\mathbb{Z})$ be arbitrary. Then there is an open normal subgroup $N\triangleleft G$ and a $d\in H^{n}(G/N,\mathbb{Z}/p\mathbb{Z})$ such that $c=\pi^{\bullet}_{N}(d)$. Let $\phi_{N}\colon\underline{H}^{n}_{\mathfrak{A}}(G/N,\mathbb{Z}/p\mathbb{Z})\to\underline{H}^{n}_{\mathfrak{A}}(G,N,\mathbb{Z}/p\mathbb{Z})$ be the map induced by the inclusion functor $\mathfrak{A}(G,N)\to\mathfrak{A}(G/N)$. Clearly $q_{G}(c)=\mathfrak{a}_{N}(\phi_{N}(q_{G/N}(d)))$, and hence the claim holds. ∎
@@
$q_{G}\colon H^{\bullet}(G,\mathbb{Z}/p\mathbb{Z})\to H^{\bullet}_{\mathfrak{A}}(G,\mathbb{Z}/p\mathbb{Z})$
@@
This theorem was proved by Quillen when $G$ is finite (see [22, Theorem 7.1 on page 567] which is actually a more general result). At the end of [20] (see 8.6 and 8.7 on pages 279–80) Scheiderer pointed out that there is an easy limit argument to derive the theorem above as a corollary. In fact, he proved that $q_{G}$ satisfies property $(i)$ in Definition 7.12. We complete his argument by showing property $(ii)$ for the convenience of the reader.
Notation 7.22.
For every pair $M,N$ of open normal subgroups of $G$ such that $M\subseteq N$, let $\pi_{M,N}\colon G/M\to G/N$ be the quotient map, and let $\mathfrak{A}(G,M,N)$ denote the image of the functor $\mathfrak{A}(\pi_{M,N})\colon\mathfrak{A}(G/M)\to\mathfrak{A}(G/N)$.
@@
For every open normal subgroup $N\triangleleft G$, there is open normal subgroup $M\triangleleft G$ such that $M\subseteq N$ and $\mathfrak{A}(G,M,N)=\mathfrak{A}(G,N)$.